OpenStudy (anonymous):

Write the solution in lowest terms. -1/3 + 3/4 = A. 5/12 B. 2/12 C. 1/6 Write the solution in lowest terms. -1/3 + 3/4 = A. 5/12 B. 2/12 C. 1/6 @Mathematics

OpenStudy (anonymous):

A

OpenStudy (anonymous):

tell me how you got it please?

OpenStudy (anonymous):

-1/3= -4/12 3/4= 9/12 -4/12+9/12=5/12

OpenStudy (anonymous):

get a common denominator first (3 times 4 = 12)

OpenStudy (anonymous):

then multiply the numerators by what you did to the denominators (so -1 times 4 = -4) and (3 times 3 = 9)

OpenStudy (anonymous):

so you have -4/12 and 9/12... add them... which equals 5/12

OpenStudy (anonymous):

get the explanation?

OpenStudy (turingtest):

\[{-1\over3} + {3\over4}={-1\over3}*{4\over4} + {3\over4}*{3\over3}={-4\over12}+{9\over12}={5\over12}\]